40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 577     Entry time: Thu Jun 26 18:29:44 2008
Author: Jenne 
Type: Update 
Category:  
Subject: MC Back online 
Jenne, Rob, Yoichi

I was playing with the Mode Cleaner earlier today, working on measuring the effect of the new filter (see next post for loop gain measurements etc.), and bumped something which made it so the Mode Cleaner would not lock.

After much poking around by Rob and Yoichi, we found that the TNC-to-2 pin LEMO from Out1 of the MC Servo Board to Ch1 of the Pentek Generic board to the right of the MC Servo Board was bad. If we touched the cable, the MC would lock, but as soon as we let go, the MC would fall out of lock. The LEMO end of the cable was not heat shrink wrapped, so the 2 wires could have been touching. I replaced the cable, and the MC locked immediately after plugging it in, so I think that has fixed the problem.
ELOG V3.1.3-